OpenStudy (anonymous):

Ann and Tom want to establish a fund for their grandsons college education. What lump sum must they deposit at a 6.9% annual interest rate, compounded quarterly, to have 40,000 in the fund at the end of 10 years

OpenStudy (kropot72):

\[A=P(1+\frac{r}{4})^{t}\] A = the amount after t quarters = 40,000 P = the lump sum originally invested r = annual interest rate as a decimal = 0.069 t = time in quarters = 10 * 4 = 40 Now just substitute in the equation and solve for P.

OpenStudy (anonymous):

Got it. Ty :)

OpenStudy (anonymous):

I thought I had it but did something wrong. Could you possibly help a little more?

OpenStudy (kropot72):

\[40000=P(1+\frac{0.069}{4})^{40}\] \[P=\frac{40000}{1.01725^{40}}\]

OpenStudy (anonymous):

Thanks so much!!

OpenStudy (kropot72):

You're welcome :)